OpenStudy (anonymous):

Simplify the expression: \[\left( \frac{ x ^{4}y ^{8} }{ x ^{6}y ^{6}} \right)^{1/2}\] Write your answer without negative exponents.

OpenStudy (anonymous):

I'm just gonna be honest here... I look at this expression and cringe in fear. So I'm gonna need help from the ground up.

OpenStudy (anonymous):

you can start by dividing each exponent by 2, since everything is being raised to the power of one half

OpenStudy (anonymous):

i.e start with \[\frac{x^2y^4}{x^3y^3}\]

OpenStudy (anonymous):

then subtract the exponents of the like terms

OpenStudy (anonymous):

Okay, then we add the exponents with the same bases?

OpenStudy (anonymous):

no, you subtract because it is a fraction (division) if it was multiplication you would add them

OpenStudy (anonymous):

Okay, so we get \[x ^{-1}y\]

OpenStudy (anonymous):

since you have to write with positive exponents, subtract the smaller ones from the bigger ones

OpenStudy (anonymous):

yeah that is right, what you wrote

OpenStudy (anonymous):

but as i said, you want positive exponents, so i would think like this \[\large \frac{x^2y^4}{x^3y^3}=\frac{y^{4-3}}{x^{3-2}}=\frac{y}{x}\]

OpenStudy (anonymous):

you got the same answer

OpenStudy (anonymous):

so its just xy?

OpenStudy (anonymous):

no it is what i wrote

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

writing with positive exponents does not mean change the minus signs to plus signs it means use the fact that \(x^{-1}y=\frac{y}{x}\)

OpenStudy (anonymous):

notice you could have done the subtractions first, like this \[\left( \frac{ x ^{4}y ^{8} }{ x ^{6}y ^{6}} \right)^{1/2}=\left( \frac{ y ^{2} }{ x ^{2}} \right)^{1/2}\] then divided the exponents by 2 either way works,whatever you think is easiest